OpenStudy (anonymous):

Differentiate (sqrt(x^9)/sqrt(0.01x)

OpenStudy (unklerhaukus):

\[0.1^2=0.01\]

OpenStudy (lgbasallote):

uhmm \[\frac{d}{dx} \left (\frac{\sqrt{x^9}}{\sqrt{0.01x}} \right)\] is that the question?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

sounds like a quotient rule question

OpenStudy (anonymous):

no, the quotient rule is not in our syllabus. ive got to\[= x^4\div \sqrt{0.01}\]

OpenStudy (lgbasallote):

you're right... \[\frac{\sqrt{x^9}}{\sqrt{0.01x}} \implies \frac{x^4\sqrt{x}}{\sqrt{0.01x}} \implies \frac{x^4}{\sqrt{0.01}}\]

OpenStudy (anonymous):

but i need to get 40x^3

OpenStudy (lgbasallote):

factor out the constant \[\frac{d}{dx} \left(\frac{x^4}{\sqrt{0.01}}\right) \implies \frac{1}{\sqrt{0.01}} \frac{d}{dx} \left(x^4\right)\]

OpenStudy (lgbasallote):

\[\large \implies \frac{4x^3}{\sqrt{0.01}}\] agree?

OpenStudy (anonymous):

ah yes, i see, ive got it now

OpenStudy (unklerhaukus):

\[0.1^2=0.01\] so \[\sqrt{0.01}=0.1\]

OpenStudy (lgbasallote):

wonderful!

OpenStudy (anonymous):

Thanks fellas!

OpenStudy (lgbasallote):

welcome ^_^ another day of saving the day for the avatar